# Foundational GPT Model for MEG

### v1

@@ 3.1 Multi-channel Wavenet
Here we describe how we adapted the Wavenet architecture (van den Oord et al.,, 2016) for electrophysiological data. Wavenet models the conditional probability of each time sample given all preceding samples autoregressively:
@@
where $\mathbf{x}_{t}$ is the sample at time $t$ and $T$ is the total sequence length. The network predicts a categorical distribution over tokenised samples using a softmax output layer. Throughout this paper we use tokenisation and quantisation interchangeably. Both have the aim of discretising a continuous quantity into a finite set of distinct bins/levels/tokens.
In the original paper, the audio waveform is tokenised using a quantisation to 8 bits following a $\mu$-law companding transform (Lewis and MTSA,, 1997):
@@
where $\mu$ controls the number of quantisation levels, set to 255 as in the original Wavenet. $f(.)$ is applied to each value of $\mathbf{x}_{t}$ independently. This nonlinear transformation improves reconstruction versus uniform quantisation of the raw input, as it skews the distribution such that more levels are allocated to smaller magnitudes. For MEG data, we observe similar benefits when applying this transform prior to quantisation. Note that the input must be scaled to $(-1,1)$ first, and clipping outliers above some threshold helps ensure a more uniform mapping.
Critically, tokenisation, in this case through quantisation, enables modelling of probability distributions over data and sampling, instead of just point estimates from MSE-based training. Cross-entropy loss also avoids the mean-prediction bias induced by MSE (Banville et al.,, 2021).
@@
In WavenetFullChannel, we first apply an embedding layer to the tokenised data, learned separately per channel. The embedding layer represents each discrete bin as a high-dimensional continuous vector, enabling powerful representations in the convolutional layers whose input channels match the embedding size. To be clear in this univariate approach the same model is applied to each channel. However, a different embedding layer is learned for each channel, meaning that for example the quantised value of 0.42 in channel x will have a different vector representation than in channel y. This helps the model differentiate between channels.
@@
where $\mathbf{Y}\in\mathbb{R}^{T\times N}$ contains the condition index $n\in(1,\dots,N)$ at each time point, and $\mathbf{O}\in\mathbb{R}^{T\times S}$ contains the subject index $s\in(1,\dots,S)$ at each time point $t\in(1,\dots,T)$. $\mathbf{W}_{y}\in\mathbb{R}^{N\times E_{n}}$ and $\mathbf{W}_{o}\in\mathbb{R}^{S\times E_{s}}$ are embedding matrices mapping the labels to learned continuous vectors of size $E_{n}$ and $E_{s}$, respectively. The subject index is the same across time points of the recording from the same subject. The condition index is set to the (visual) stimuli presented (e.g., one of the 118 images in Cichy et al., (2016)), for exactly those time points when the stimulus is on. At any other time, the task condition embedding $\mathbf{H}_{y}$ is set to 0.

### v2

@@ A.1 Multi-channel Wavenet
Here we describe how we adapted the Wavenet architecture (van den Oord et al., 2016) for electrophysiological data. Wavenet models the conditional probability of each time sample given all preceding samples autoregressively:
@@
where $\mathbf{x}_{t}$ is the sample at time $t$ and $T$ is the total sequence length. Throughout this paper we use tokenisation and quantisation interchangeably. Both have the aim of discretising a continuous quantity into a finite set of distinct bins/levels/tokens.
In the original paper, the audio waveform is tokenised using a quantisation to 8 bits following a $\mu$-law companding transform (Brokish and Lewis, 1997):
@@
where $\mu$ controls the number of quantisation levels, set to 255 as in the original Wavenet. $f(.)$ is applied to each value of $\mathbf{x}_{t}$ independently. This nonlinear transformation improves reconstruction versus uniform quantisation of the raw input, as it skews the distribution such that more levels are allocated to smaller magnitudes. For MEG data, we observe similar benefits when applying this transform prior to quantisation. Note that the input must be scaled to (1,1)(-1,1) first, and clipping outliers above some threshold helps ensure a more uniform mapping.
@@
In WavenetFullChannel, we first apply an embedding layer to the tokenised data, learned separately per channel. To be clear in this univariate approach the same model is applied to each channel. However, a different embedding layer is learned for each channel, meaning that for example the quantised value of 0.42 in channel x will have a different vector representation than in channel y. This helps the model differentiate between channels.
@@
where $\mathbf{Y}\in\mathbb{R}^{T\times N}$ contains the condition index $n\in(1,\dots,N)$ at each time point, and $\mathbf{O}\in\mathbb{R}^{T\times S}$ contains the subject index $s\in(1,\dots,S)$ at each time point $t\in(1,\dots,T)$. $\mathbf{W}_{y}\in\mathbb{R}^{N\times E_{n}}$ and $\mathbf{W}_{o}\in\mathbb{R}^{S\times E_{s}}$ are embedding matrices mapping the labels to learned continuous vectors of size $E_{n}$ and $E_{s}$, respectively. The subject index is the same across time points of the recording from the same subject. The condition index is set to the (visual) stimuli presented (e.g., one of the 118 images in Cichy et al. (2016)), for exactly those time points when the stimulus is on. At any other time, the task condition embedding $\mathbf{H}_{y}$ is set to 0.